# TE-NeXt: A LiDAR-Based 3D Sparse Convolutional Network for Traversability Estimation

### v3

@@ Abstract
Among the multitude of scenarios in which autonomous robots are intended to operate, natural environments present the most significant challenges in the context of traversability estimation. Conversely, although comparatively less complex, structured environments are the most prevalent in the broader sense of autonomous navigation due to the abundance of human-designed spaces, which offer more predictable features for navigation systems. Therefore, both of these considerations lead to the conclusion that there is a need to develop methods that are specifically tailored to unstructured environments, while ensuring that the performance in structured environments is not significantly compromised. Existing methods are vulnerable to the influence of stationary environmental conditions or are constrained by the platform or vehicle on which they are developed. In this context, this paper makes a contribution to the state of the art with TE-NeXt. This network is a customised and efficient architecture for traversability estimation from sparse LiDAR point clouds based on a encoder-decoder topology that includes several modifications regarding: (i) the input features; (ii) the structure of encoder-decoder resolution levels; and (iii) the constitution of the 3D convolutional block. The modifications provide the robot with a deeper understanding of traversability in natural and unstructured environments. Thus, the designed architecture presents a trade-off solution, exhibiting superior performance in unstructured environments (the Rellis-3D dataset) while maintaining high reliability and robustness in urban environments (the SemanticKITTI dataset). The implementation of the presented approach is available at an open repository to ensure result reproducibility.
@@
journal: Expert Systems with Applications
@@ 1 Introduction
To enhance the efficacy of specialised techniques for off-road traversability estimation, this study proposes TE-NeXt, a LiDAR-based 3D sparse convolutional architecture capable of discerning traversable and non-traversable areas captured by a LiDAR sensor, enriching the perception of the environment with semantic information in addition to geometric information. In summary, the main contributions of this paper are:
@@
    The applicability of the proposed method in an autonomous navigation system in dynamic environments.
@@ 2.1 Vision-based traversability estimation
One of the earliest supervised methods was proposed in [13], where a Multi-Layer Perceptron (MLP) classifier combined with colour and texture features (extracted from RGB images) was used as a method for automatic classification of road and off-road terrain. However, the methods rapidly evolved towards the utilisation of visual information as an input, largely due to the impressive results demonstrated by convolutional networks in image classification tasks. In this context, network architectures such as GA-Nav [14] and DeconvNet [15] emerged. Both solely employ visual information as an input for terrain classification, with the objective of identifying safe traversal routes. Additionally, both approaches employed semantic segmentation, grouping the results into six and five classes, respectively.
In addition to methods that use visible spectrum images, methods have also been developed that rely on the infrared spectrum to estimate the traversability of the environment in order to avoid low-light situations such as night, backlight, haze, etc… Pioneering works such as [16] present a convolutional neural network (CNN) for traversability estimation, and compare the optimal way to implement multi-model data fusion, now known as early or late fusion. URTSegNet [17], inspired by methods that perform segmentation in infrared images [18], proposes a real-time neural network including two diferent branches to extract spatial detail and contextual features.
@@ 2.2 Geometric and hybrid methods for traversability estimation
A common practice in a three dimensional context is 3D voxel partition, in approaches such as [23, 24] the authors represent the point cloud as a 3D voxel occupancy map. This voxel representation avoids the need for commonly used elevation maps, which are susceptible to inaccuracies in the presence of overhanging obstacles and in multi-storey or low ceiling scenarios. In addition to the internal representation of the environment, they proposed a 3D sparse encoder-decoder that exploits the sparsity of the voxel representation for efficient computation. However, there are various methods of representing point clouds. An approach that is outlined in [25], involves the arrangement of point clouds within a multilevel polar grid, which naturally takes into account the non-uniform density of point clouds. An alternative approach is presented in [26], where point clouds are represented using cylindrical voxels based on the cylindrical coordinate system.
As with 2D vision, methods that fused input information of different nature were developed in the 3D context. One of the most common proposals is the merging of surface normals extracted from LiDAR data with RGB images [27, 28, 29]. All these studies utilise multi-modal fusion of RGB images and geometric information derived from LiDAR point clouds. The proposed architectures employ transformer-based encoder blocks to extract features from the input modalities.
@@ 2.4 Current proposal
This paper presents TE-NeXt, an encoder-decoder architecture optimised for traversability estimation in unstructured environments. The approach presented here can be classified as a supervised method, in accordance with the paradigm described in Subsection 2.2, as it is a supervised paradigm method. However, it employs local point clouds, a distinctive feature that sets it apart from other methods, such as those described in [23, 24], which utilise point cloud maps and preprocess the data. Moreover, in contrast to the techniques outlined in Subsection 2.1, TE-NeXt employs an optical sensor that is invariant to changes in light conditions (LiDAR sensor). Furthermore, the proposed methodology is designed to overcome the limitations associated with self-supervised learning approaches. Unlike existing methods, it does not require a specific platform, making it a more general and secure solution for all types of vehicles. Table 1 outlines the distinctive features of the proposed approach in comparison to existing, state-of-the-art methodologies.
@@ 3.2 TE-NeXt residual block
The N-channel input first passes through a 1x1x1 convolutional layer that reduces the channels by a factor of 4, aleviating the computational complexity and memory requirements of subsequent layers. It operates on each channel independently, permitting learning channel-specific filters and weights. This allows the network to selectively suppress redundant information across channels. It then passes through a 3D convolutional layer with a 7x7x7 kernel, maintaining the input dimensions of the layer.
@@ 4.3.2 Global Design
Steps ID Experimental Modifications
@@ 4.3.3 Micro Design
  • GeLu activation function. The ReLU has been the most popular activation function in deep learning due to its simplicity and efficiency. However, recent Transformer models have started using the GELU instead. GELU is a smoother variant of ReLU that can enable better training of deep networks by reducing issues like vanishing gradients and allowing useful gradients for negative input values. This activation function, experiment M.6, has resulted in an enhancement in performance relative to its predecessor, with an F1 metric of 80.5%.
@@ 5 Results
Table 4 presents the results of several semantic segmentation methods for the estimation of traversability in urban environments, using the validation sequence of the SemanticKITTI dataset.
@@ 5.5 Implementation details
Table 6 provides an informative comparison of key efficiency metrics for the methods compared above. The results highlight the substantial differences in model size, memory needs and latency that can arise from architecture choices. With regard to the model size, it is evident that TE-NeXt fulfils one of the primary objectives of its design: the reduction of the model’s complexity without compromising its performance.
Methods using 3D convolutions (MinkU-Net34C, MinkU-Net+ConvNeXt, TE-NeXt) tend to have lower training memory requirements compared to those using MLP or 2D convolutions, despite having more parameters in some cases. This suggests 3D convolutions may be more memory efficient during training.
Method’s Efficiency Params
Type
Training
Memory
Inference
Latency
PointNet [54] 3.5M MLP 13GB 11 ms
PointNet++ [55] 0.9M MLP 13.6GB 250 ms
SqueezeSegv3 [56] 9.2M 2D 37.8GB 29 ms
SalsaNext [51] 6.7M 2D 11.2GB 28 ms
MinkU-Net [39] 37.9M 3D 6.1GB 53ms
MinkU-Net [39]+ConvNeXt [10] 66.4M 3D 8.7GB 78 ms
TE-NeXt (ours) 5.4M 3D 5.4GB 31 ms
@@ 6 Conclusion
Furthermore, the comparison results presented in this manuscript validate the effectiveness of the proposed approach, TE-NeXt, for traversability estimation in highly natural environments, since the network outperform state-of-the-art methods achieving the best results in terms of model stability (F1 score), mIOU and TPR. Nevertheless, further research and improvements to existing methods are required to cope with the increasing complexity of natural scenes. Additionally, the network demonstrates comparable results to the most advanced semantic segmentation methods in urban environments.

### v4

@@ Abstract
Among the multitude of scenarios in which autonomous robots are intended to operate, natural environments present the most significant challenges in the context of traversability estimation. Conversely, although comparatively less complex, structured environments are the most prevalent in the broader sense of autonomous navigation due to the abundance of human-designed spaces, which offer more predictable features for navigation systems. Therefore, both of these considerations lead to the conclusion that there is a need to develop methods that are specifically tailored to unstructured environments, while ensuring that the performance in structured environments is not significantly compromised. Existing methods are vulnerable to the influence of stationary environmental conditions or are constrained by the platform or vehicle on which they are developed. In this context, this paper makes a contribution to the state of the art with TE-NeXt (Traversability Estimation ConvNeXt). This network is a customised and efficient architecture for traversability estimation from sparse LiDAR point clouds based on a encoder-decoder topology that includes several modifications regarding: (i) the input features; (ii) the structure of encoder-decoder resolution levels; and (iii) the constitution of the 3D convolutional block. The modifications provide the robot with a deeper understanding of traversability in natural and unstructured environments. Thus, the designed architecture presents a trade-off solution, exhibiting superior performance in challenging unstructured environments (the Rellis-3D dataset), with an 82% F1 score, which outperforms existing methodologies, while maintaining high reliability and robustness in urban environments (the SemanticKITTI dataset). Furthermore, a fully autonomous navigation framework is presented in detail. The implementation of the presented approach is available at an open repository to ensure result reproducibility.
@@
journal: Results in Engineering
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
TE-NeXt: Sparse CNN for traversability estimation in complex environments.
Novel sparse convolution block inspired by ConvNeXt, surpassing ResNet Blocks.
F1-5% improvement in traversability estimation for unstructured terrain.
TE-NeXt enables autonomous navigation in dynamic environments.
@@ 1 Introduction
To enhance the efficacy of specialised techniques for off-road traversability estimation, this study proposes TE-NeXt (Traversability Estimation ConvNeXt), a LiDAR-based 3D sparse convolutional architecture capable of discerning traversable and non-traversable areas captured by a LiDAR sensor, enriching the perception of the environment with semantic information in addition to geometric information. In summary, the main contributions of this paper are:
@@
    A real-world application of the proposed method in an autonomous navigation system in a semi-structured environment.
@@ 2.1 Vision-based traversability estimation
One of the earliest supervised methods was proposed in [13], where a Multi-Layer Perceptron (MLP) classifier combined with colour and texture features (extracted from RGB images) was used as a method for automatic classification of road and off-road terrain. However, the methods rapidly evolved towards the utilisation of visual information as an input, largely due to the impressive results demonstrated by convolutional networks in image classification tasks. In this context, network architectures such as GA-Nav [14] and DeconvNet [15] emerged. Both solely employ visual information as an input for terrain classification, with the objective of identifying safe traversal routes. Additionally, both approaches employed semantic segmentation, grouping the results into six and five classes, respectively. The reliability of these techniques is fundamentally limited by environmental dynamics and data capture constraints during vehicle-based image acquisition [16].
In addition to methods that use visible spectrum images, methods have also been developed that rely on the infrared spectrum to estimate the traversability of the environment in order to avoid low-light situations such as night, backlight, haze, etc… Pioneering works such as [17] present a convolutional neural network (CNN) for traversability estimation, and compare the optimal way to implement multi-model data fusion, now known as early or late fusion. URTSegNet [18], inspired by methods that perform segmentation in infrared images [19], proposes a real-time neural network including two diferent branches to extract spatial detail and contextual features. Infrared imaging is significantly less effective in scenarios characterized by low thermal contrast between environmental elements. As a result, sophisticated and highly specialised calibration methods are required to derive precise measurements [20].
@@ 2.2 Geometric and hybrid methods for traversability estimation
A common practice in a three dimensional context is 3D voxel partition, in approaches such as [25, 26] the authors represent the point cloud as a 3D voxel occupancy map. This voxel representation avoids the need for commonly used elevation maps, which are susceptible to inaccuracies in the presence of overhanging obstacles and in multi-storey or low ceiling scenarios. In addition to the internal representation of the environment, they proposed a 3D sparse encoder-decoder that exploits the sparsity of the voxel representation for efficient computation. However, there are various methods of representing point clouds. An approach that is outlined in [27], involves the arrangement of point clouds within a multilevel polar grid, which naturally takes into account the non-uniform density of point clouds. An alternative approach is presented in [28], where point clouds are represented using cylindrical voxels based on the cylindrical coordinate system. These approaches mitigate the natural dispersion of data by creating three-dimensional maps in advance, which results in higher computational costs.
As with 2D vision, methods that fused input information of different nature were developed in the 3D context. One of the most common proposals is the merging of surface normals extracted from LiDAR data with RGB images [29, 30, 31]. All these studies utilise multi-modal fusion of RGB images and geometric information derived from LiDAR point clouds. The main issue associated with the integration of different data types may be attributed to the misalignment of the timelines of the two sensors, which can lead to projection errors and generate inaccurate information.
Table 1: Overview of related works.
@@ 2.4 Current proposal
This paper presents TE-NeXt, an encoder-decoder architecture optimised for traversability estimation in unstructured environments. The approach presented here can be classified as a supervised method, in accordance with the paradigm described in Subsection 2.2, as it is a supervised paradigm method. However, it employs local point clouds, a distinctive feature that sets it apart from other methods, such as those described in [25, 26], which utilise point cloud maps and preprocess the data. Moreover, in contrast to the techniques outlined in Subsection 2.1, TE-NeXt employs an optical sensor that is invariant to changes in light conditions (LiDAR sensor). Furthermore, the proposed methodology is designed to overcome the limitations associated with self-supervised learning approaches. Unlike most of the existing methods, it does not require a specific platform, making it a more general and secure solution for all types of vehicles. Table 1 outlines the distinctive features of the proposed approach in comparison to existing, state-of-the-art methodologies.
@@ 3 TE-NeXt: AN ARCHITECTURE FOR TRAVERSABILITY ESTIMATION
3.1 Mathematical Formulation
The traversability estimation problem is formulated as a semantic segmentation task, where a point cloud $\mathcal{P}$ with $N$ points is characterized by Cartesian coordinates $\vec{p_{i}}=(x_{i},y_{i},z_{i})\in\mathbb{R}^{3}$ and feature vectors $\vec{f_{i}}\in\mathbb{R}^{d_{\text{in}}}$. The objective is to learn a mapping function $f:\mathbb{R}^{3+d_{\text{in}}}\rightarrow[0,1]$ that assigns a traversability probability $\hat{l}_{i}$ to each point, where $\hat{l}_{i}\in[0,1]$. By training on a labeled dataset, $\{(\vec{p}_{i},\vec{f}_{i}),\hat{l}_{i}\}$, the model aims to capture intricate spatial and feature-based patterns, enabling precise prediction of point-wise traversability through a deep learning approach that minimizes classification error across the entire point cloud representation.
@@ 3.3 TE-NeXt residual block
The N-channel input first passes through a 1x1x1 convolutional layer that reduces the channels by a factor of 4, aleviating the computational complexity and memory requirements of subsequent layers. The idea behind this operation is to facilitate the independent examination of cross-channel correlations and spatial correlations. Therefore, it operates on each channel independently, allowing channel-specific filters and weights to be learned. The 1x1x1 convolution provides a sophisticated mechanism for neural networks to simultaneously reduce computational complexity and mitigate inherent spatial and channel redundancies prevalent in standard convolutional architectures, enabling more efficient feature representation through intelligent channel-wise transformations and selective information preservation [44]. The underlying premise is that cross-channel and spatial correlations are sufficiently decoupled that mapping them together is not the optimal approach [45]. Once the aforementioned mapping has been completed, the spatial correlations are mapped using regular 3D convolutions with a 7x7x7 kernel.
@@ 4.3 Ablation study
To facilitate the comprehension of the ablation study, Table 3 has been introduced as a legend, providing information on the starting point of each experiment and the changes made. Additionally, the number of parameters of each model is included, revealing a notable reduction in the number of parameters compared to the starting point. A closer examination reveals that the initial block of experiments has no discernible effect on the network’s trainable parameters. However, in subsequent blocks of experiments, notable alterations in the model size emerge. These changes are due to modifications in the convolution operations, kernel sizes, and the number of parallel branches in the convolutional blocks.
@@ 4.3.2 Global Design
Steps ID Experimental Modifications Params
@@ 4.3.3 Micro Design
  • GeLu activation function. The ReLU has been the most popular activation function in deep learning due to its simplicity and efficiency. However, recent Transformer models have started using the GELU instead [54, 10, 55]. GELU is a smoother variant of ReLU that can enable better training of deep networks by reducing issues like vanishing gradients and allowing useful gradients for negative input values. This activation function, experiment M.6, has resulted in an enhancement in performance relative to its predecessor, with an F1 metric of 80.5%.
@@ 5 Results and Discussions
5.1 Evaluation metrics
The proposed method and other competitive state-of-the-art methods are evaluated based on the fundamental components of the confusion matrix: true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN). Table 4 presents the equations for each calculated metric. Specifically, the accuracy metric measures the percentage of correctly classified instances and it is highly susceptible to bias due to the unequal distribution of instances across classes in an unbalanced dataset. In contrast, the F1 score metric provides a more reliable assessment as it is the harmonic mean between precision and recall, which aim to minimise false positives and false negatives, respectively. The Intersection over Union (IoU) is specifically computed as the ratio of the intersection area to the union area between the predicted and ground truth segmentation labels. The mean IoU (mIoU) is then computed by averaging the IoU values across all classes, ensuring each class contributes equally to the final performance metric. Finally, the true positive and true negative rate demonstrates the proportion of correctly classified true positive and true negative instances.
Table 4: Performance metrics formulation.
Metric Definition
Accuracy $\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$
F1 score $2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
mIoU $\frac{1}{2}\left(\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}+\frac{\text{TN}}{\text{TN}+\text{FP}+\text{FN}}\right)$
True Positive Rate $\frac{\text{TP}}{\text{TP}+\text{FN}}$
True Negative Rate $\frac{\text{TN}}{\text{TN}+\text{FP}}$
@@ 5.3 Performance in natural environments
In order to complete the comparison, Table 7 provides a detailed quantitative analysis of the aforementioned methods in terms of efficiency. The results highlight significant differences in model size, memory requirements, and latency that can arise from architecture choices. Regarding to the model size, it is evident that TE-NeXt achieves one of its primary design goals: the reduction of the model’s complexity without compromising its performance. Methods using 3D convolutions (MinkU-Net34C, MinkU-Net+ConvNeXt, TE-NeXt) tend to have lower training memory requirements compared to those using MLP or 2D convolutions, despite having more parameters in some cases. This suggests 3D convolutions may be more memory efficient during training.
Method’s Efficiency Params
Type
Training
Memory
Inference
Latency
PointNet [60] 3.5M MLP 13GB 11 ms
PointNet++ [61] 0.9M MLP 13.6GB 250 ms
SqueezeSegv3 [62] 9.2M 2D 37.8GB 29 ms
SalsaNext [57] 6.7M 2D 11.2GB 28 ms
MinkU-Net [41] 37.9M 3D 6.1GB 53ms
MinkU-Net [41]+ConvNeXt [10] 66.4M 3D 8.7GB 78 ms
TE-NeXt (ours) 5.4M 3D 5.4GB 31 ms
@@ 6 Conclusion
Furthermore, the comparison results presented in this manuscript validate the effectiveness of the proposed approach, TE-NeXt, for traversability estimation in highly natural environments, achieving results of 82% on the F1 metric, composed of the harmonic mean of precision and recall, and 77% on the mIoU metric, among others. It thus demonstrates superior performance compared to the current state-of-the-art methods, offering enhanced stability in terms of performance. Nevertheless, further research and improvements to existing methods are required to cope with the increasing complexity of natural scenes. Additionally, the network demonstrates comparable results to the most advanced semantic segmentation methods in urban environments.